OpenStudy (anonymous):

Help! I dont understand this question: A quadratic function has a vertex at (2,5) and passes through the point (4,9). What is the correct algorithm for finding an equation for the quadratic function?

OpenStudy (anonymous):

A) y = a (x - 2)^2 + 5 9 = a(4-2)^2 + 5 a = 1 y = (x - 2)^2 + 5 B) y = a (x - 4)^2 + 9 5 = a(2 - 4)^2 + 9 a = -1 y = -1(x - 4)2 + 9 C) y = a (x - 5)^2 + 2 9 = a(4-5)^2 + 2 a = 7 y = 7(x - 5)2 + 2 D) y = a (x - 9)^2 + 4 5 = a(2 - 9)^2 + 4 a = 1/49 y = 1/49(x - 4)2 + 9

ganeshie8 (ganeshie8):

vertex form of quadratic equation : \(\large y = a(x-h)^2 + k\) where : vertex = \((h,k)\)

OpenStudy (anonymous):

oh ok . . . um is h: 5 and k: 9

ganeshie8 (ganeshie8):

Nope.

ganeshie8 (ganeshie8):

A quadratic function has a vertex at (2, 5) h k

ganeshie8 (ganeshie8):

^^

ganeshie8 (ganeshie8):

h : 2 k : 5

ganeshie8 (ganeshie8):

plug them in the quadratic : \(\large y = a(x-2)^2 + 5 \)

ganeshie8 (ganeshie8):

you still need to find the value of \(a\)

hero (hero):

Just to make it clearer, to add to what @ganshie8 said, The vertex form of a parabola is \(\large y = a(x-h)^2 + k\) where \((h,k)\) is the vertex and \((x,y)\) is one of the points that goes through its graph.

hero (hero):

In this case you were given \((h,k) = ((2,5)\) and \((x,y) = (4,9)\) Once you've inserted both the vertex and the point into the vertex formula, finding the value of \(a\) should be effortless.

hero (hero):

@ganeshie8 :)

OpenStudy (anonymous):

Alright got it thanks :D